# Quantum stochastic thermodynamics in the mesoscopic-leads formulation

### v1

@@ C.3 Energy current
$\begin{split}I^{r}_{E_{\alpha}}(t)\mathrm{d}t=&\ \mathrm{Tr}\left[B_{\alpha}\left(C^{r}HC^{r}-\frac{1}{2}(HC^{r}+C^{r}H)\right)\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[H\bar{C}^{r}\left(1-\Lambda^{+}_{\alpha}\right)F_{\alpha}\bar{C}^{r}\right]\mathrm{d}t\\ &-\mathrm{Tr}\left[HC^{r}\left(1-\Lambda^{-}_{\alpha}\right)(\Gamma_{\alpha}-F_{\alpha})C^{r}\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[H\bar{C}^{r}\left(\frac{\Lambda^{+}\mathrm{d}N^{+}_{\alpha}(t)}{\bar{C}}\right)\bar{C}^{r}\right]\\ &-\mathrm{Tr}\left[HC^{r}\left(\frac{\Lambda^{-}\mathrm{d}N^{-}_{\alpha}(t)}{C^{r}}\right)C^{r}\right].\end{split}$ (119)
@@ C.4 Measurement energy current
$\begin{split}I_{E_{\mathrm{M}}}\mathrm{d}t=&-\frac{\mathrm{d}t}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\mathrm{Tr}\left[H\{{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k},\rho_{r}\}\right]+\mathrm{d}t\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\langle H\rangle_{r}\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\Lambda_{k}^{\sigma}\mathrm{d}N^{\sigma}_{k}\frac{\mathrm{Tr}\left[{L^{\sigma}_{k}}^{\dagger}H_{\mathrm{int}}L^{\sigma}_{k}\rho_{r}\right]}{\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}}+\mathrm{d}t\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\mathrm{Tr}\left[{L^{\sigma}_{k}}^{\dagger}H_{\mathrm{int}}L^{\sigma}_{k}\rho_{r}\right]\\ &+\frac{1}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\Lambda_{k}^{\sigma}\mathrm{d}N^{\sigma}_{k}\frac{\mathrm{Tr}\left[\{{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k},H_{0}\}\rho_{r}\right]}{\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}}-\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\Lambda_{k}^{\sigma}\mathrm{d}N^{\sigma}_{k}\langle H\rangle_{r}\\ &+\frac{\mathrm{d}t}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\mathrm{Tr}\left[H_{0}\{{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k},\rho_{r}\}\right]-\mathrm{d}t\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\langle H\rangle_{r}\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}.\end{split}$ (120)
@@
$\begin{split}I_{E_{\mathrm{M},\alpha}}\mathrm{d}t=&\mathrm{Tr}\left[B_{\alpha}\left(C^{r}HC^{r}-\frac{1}{2}(HC^{r}+C^{r}H)\right)\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[\left(1-\Lambda^{+}_{\alpha}\right)F_{\alpha}\bar{C}^{r}H_{\mathrm{int}}\bar{C}^{r}\right]\mathrm{d}t-\mathrm{Tr}\left[\left(1-\Lambda^{-}_{\alpha}\right)(\Gamma_{\alpha}-F_{\alpha})C^{r}H_{\mathrm{int}}C^{r}\right]\mathrm{d}t\\ &-\frac{1}{2}\mathrm{Tr}\left[\left(1-\Lambda^{+}_{\alpha}\right)F_{\alpha}\left(\bar{C}^{r}H_{0}C^{r}+C^{r}H_{0}\bar{C}^{r}\right)\right]\mathrm{d}t-\frac{1}{2}\mathrm{Tr}\left[\left(1-\Lambda^{-}_{\alpha}\right)(\Gamma_{\alpha}-F_{\alpha})\left(\bar{C}^{r}H_{0}C^{r}+C^{r}H_{0}\bar{C}^{r}\right)\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[\left(\frac{\Lambda^{+}\mathrm{d}N^{+}_{\alpha}}{\bar{C}^{r}}\right)\bar{C}^{r}H_{\mathrm{int}}\bar{C}^{r}\right]-\mathrm{Tr}\left[\left(\frac{\Lambda^{-}\mathrm{d}N^{-}_{\alpha}}{C^{r}}\right)C^{r}H_{\mathrm{int}}C^{r}\right]\\ &-\frac{1}{2}\mathrm{Tr}\left[\left(\frac{\Lambda^{+}\mathrm{d}N^{+}_{\alpha}}{\bar{C}^{r}}\right)\left(\bar{C}^{r}H_{0}C^{r}+C^{r}H_{0}\bar{C}^{r}\right)\right]-\frac{1}{2}\mathrm{Tr}\left[\left(\frac{\Lambda^{-}\mathrm{d}N^{-}_{\alpha}}{C^{r}}\right)\left(\bar{C}^{r}H_{0}C^{r}+C^{r}H_{0}\bar{C}^{r}\right)\right].\end{split}$ (121)

### v2

@@ C.3 Energy current
$\begin{split}I^{r}_{E_{\alpha}}(t)\mathrm{d}t=&\ \mathrm{Tr}\left[B_{\alpha}\left(C^{r}HC^{r}-\frac{1}{2}(HC^{r}+C^{r}H)\right)\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[H\bar{C}^{r}\left(1-\Lambda^{+}_{\alpha}\right)F_{\alpha}\bar{C}^{r}\right]\mathrm{d}t\\ &-\mathrm{Tr}\left[HC^{r}\left(1-\Lambda^{-}_{\alpha}\right)(\Gamma_{\alpha}-F_{\alpha})C^{r}\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[H\bar{C}^{r}\left(\frac{\mathrm{d}N^{+}_{\alpha}(t)}{\bar{C}}\right)\bar{C}^{r}\right]\\ &-\mathrm{Tr}\left[HC^{r}\left(\frac{\mathrm{d}N^{-}_{\alpha}(t)}{C^{r}}\right)C^{r}\right].\end{split}$ (117)
@@ C.4 Measurement energy current
$\begin{split}I_{E_{\mathrm{M}}}\mathrm{d}t=&-\frac{\mathrm{d}t}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\mathrm{Tr}\left[H\{{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k},\rho_{r}\}\right]+\mathrm{d}t\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\langle H\rangle_{r}\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}\\ &+\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\mathrm{d}N^{\sigma}_{k}\frac{\mathrm{Tr}\left[{L^{\sigma}_{k}}^{\dagger}H_{\mathrm{int}}L^{\sigma}_{k}\rho_{r}\right]}{\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}}+\mathrm{d}t\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\mathrm{Tr}\left[{L^{\sigma}_{k}}^{\dagger}H_{\mathrm{int}}L^{\sigma}_{k}\rho_{r}\right]\\ &+\frac{1}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\mathrm{d}N^{\sigma}_{k}\frac{\mathrm{Tr}\left[\{{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k},H_{0}\}\rho_{r}\right]}{\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}}-\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}\mathrm{d}N^{\sigma}_{k}\langle H\rangle_{r}\\ &+\frac{\mathrm{d}t}{2}\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\mathrm{Tr}\left[H_{0}\{{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k},\rho_{r}\}\right]-\mathrm{d}t\sum_{k=1}^{L}\sum_{\sigma\in\{+,-\}}(1-\Lambda_{k}^{\sigma})\langle H\rangle_{r}\langle{L^{\sigma}_{k}}^{\dagger}L^{\sigma}_{k}\rangle_{r}.\end{split}$ (118)
@@
$\begin{split}I_{E_{\mathrm{M},\alpha}}\mathrm{d}t=&\mathrm{Tr}\left[B_{\alpha}\left(C^{r}HC^{r}-\frac{1}{2}(HC^{r}+C^{r}H)\right)\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[\left(1-\Lambda^{+}_{\alpha}\right)F_{\alpha}\bar{C}^{r}H_{\mathrm{int}}\bar{C}^{r}\right]\mathrm{d}t-\mathrm{Tr}\left[\left(1-\Lambda^{-}_{\alpha}\right)(\Gamma_{\alpha}-F_{\alpha})C^{r}H_{\mathrm{int}}C^{r}\right]\mathrm{d}t\\ &-\frac{1}{2}\mathrm{Tr}\left[\left(1-\Lambda^{+}_{\alpha}\right)F_{\alpha}\left(\bar{C}^{r}H_{0}C^{r}+C^{r}H_{0}\bar{C}^{r}\right)\right]\mathrm{d}t-\frac{1}{2}\mathrm{Tr}\left[\left(1-\Lambda^{-}_{\alpha}\right)(\Gamma_{\alpha}-F_{\alpha})\left(\bar{C}^{r}H_{0}C^{r}+C^{r}H_{0}\bar{C}^{r}\right)\right]\mathrm{d}t\\ &+\mathrm{Tr}\left[\left(\frac{\mathrm{d}N^{+}_{\alpha}}{\bar{C}^{r}}\right)\bar{C}^{r}H_{\mathrm{int}}\bar{C}^{r}\right]-\mathrm{Tr}\left[\left(\frac{\mathrm{d}N^{-}_{\alpha}}{C^{r}}\right)C^{r}H_{\mathrm{int}}C^{r}\right]\\ &-\frac{1}{2}\mathrm{Tr}\left[\left(\frac{\mathrm{d}N^{+}_{\alpha}}{\bar{C}^{r}}\right)\left(\bar{C}^{r}H_{0}C^{r}+C^{r}H_{0}\bar{C}^{r}\right)\right]-\frac{1}{2}\mathrm{Tr}\left[\left(\frac{\mathrm{d}N^{-}_{\alpha}}{C^{r}}\right)\left(\bar{C}^{r}H_{0}C^{r}+C^{r}H_{0}\bar{C}^{r}\right)\right].\end{split}$ (119)
@@ Appendix D Monte Carlo trajectory sampling
Appendix E Steady-state heat current and entropy production in local GKSL master equations
Figure 9: Time-evolving steady-state heat current and entropy production after the first projective measurement, shown for cases excluding (blue) and including (red) the measurement energy current. a) Time-evolving average of the integrated heat current. When the measurement energy current is excluded, a spurious heat current persists even in the steady state. Including the measurement energy current corrects this discrepancy. b) Distribution of the integrated heat current. c) Distribution of the total entropy production. The modified entropy production, $\tilde{S}_{\mathrm{tot}}$, is centered and sharply peaked around zero, as expected in a steady state. d) Convergence to the integral fluctuation theorem (Eq. (53)) for $S_{\mathrm{tot}}$, but not for $\tilde{S}_{\mathrm{tot}}$. Parameters: $L$ = 10, up to 30000 trajectories, $\omega_{\mathrm{max}}=1$, $\epsilon$ = $\omega_{\mathrm{max}}/4$, $\Gamma=\omega_{\mathrm{max}}/8$, $T=\omega_{\mathrm{max}}$, $\mu=\omega_{\mathrm{max}}/16$, $\Gamma\tau=50$.
In the following, we elaborate on why, in the mesoscopic-leads formalism or any local GKSL master equation, entropy production along trajectories remains nonzero in a steady state and why its mean is nonzero. This arises from the definition of total entropy production
$S_{\mathrm{tot}}(r_{\left[0,\tau\right]})=\log\left(\frac{p^{0}_{n}}{p^{\tau}_{m}}\right)+\Sigma_{r_{\left(0,\tau\right)}},$ (122)
where $\Sigma_{r_{\left(0,\tau\right)}}=-\frac{\Delta Q_{r}}{T}$ and
$\Delta Q_{r}=\sum_{\alpha=1}^{N_{\mathrm{R}}}\sum_{k=1}^{l_{\alpha}}\sum_{\sigma\in\{+,-\}}\int_{0}^{\tau}\frac{\Delta E^{\sigma}_{k,\alpha}-\mu_{\alpha}\Delta N^{\sigma}_{k,\alpha}}{T_{\alpha}}\mathrm{d}N_{k}(t).$ (123)
Assume the initial state is the steady state, and the first projective measurement yields outcome $n_{0}$, corresponding to the steady-state eigenstate projector $\Pi_{n_{0}}$. This state then evolves stochastically over time $\tau$. Crucially, the state at the end of this trajectory is not guaranteed to be a steady-state eigenstate. Instead, it has finite overlap with a set of steady-state eigenstates (typically of size >1>1) and may be projected into any of these by the second measurement in the two-point measurement (TPM) scheme. As a result, the relative surprisal between the two measurement outcomes cannot be uniquely associated with a specific entropy flux $\Sigma_{r_{\left(0,\tau\right)}}$. Consequently, even in a steady state, there exists a distribution of total entropy flux. The nonzero average arises because measurement energy is excluded from Eq. (52), as illustrated in Fig. 9 a)–c). Including the measurement energy yields a modified entropy production $\tilde{S}_{\mathrm{tot}}$ which, however, would violate the integral fluctuation theorem in Eq. (53Manzano and Zambrini (2022), as shown in Fig. 9 d). Interestingly, while this modified entropy production appears to violate the standard fluctuation theorem, it seems to converge rapidly to some fixed, finite value. Exploring this feature further would be an intriguing direction for future work, but it lies beyond the scope of this study. The discrepancy between $S_{\mathrm{tot}}$ and $\tilde{S}_{\mathrm{tot}}$ is expected to be suppressed with an increasing number of lead modes per bath, $l$, as the system-lead coupling, which contributes to the measurement energy current, scales as $\kappa_{k}\propto 1/\sqrt{l}$. However, due to both statistical constraints and the computational cost of determining $p_{m}^{r\tau}$ (see Sec. V), which grows exponentially with the size of the extended system even for Gaussian systems, we limited our analysis to $l=10$, given the available computational resources. Since sampling the distribution of the (integrated) energy and particle currents, as opposed to the entropy production, does not require relying on the TPM, leads with a larger number of modes become feasible, as demonstrated in Sec. VI, where $l=20$ .